# A Technical Study into 0.5B Reasoning Language Models

### v2

@@ Abstract
The ongoing evolution of language models has led to the development of large-scale architectures that demonstrate exceptional performance across a wide range of tasks. However, these models come with significant computational and energy demands, as well as potential privacy implications. In this context, Small Reasoning Language Models (SRLMs) with approximately 0.5 billion parameters present a compelling alternative due to their remarkable computational efficiency and cost-effectiveness, particularly in resource-constrained environments. Despite these advantages, the limited capacity of 0.5 billion parameter models poses challenges in handling complex tasks such as mathematical reasoning and code generation. This research investigates various training strategies—including supervised fine-tuning (SFT), knowledge distillation (KD), and reinforcement learning (RL), as well as their hybrid implementations, to enhance the performance of 0.5B SRLMs. We analyze effective methodologies to bridge the performance gap between SRLMS and larger models and present insights into optimal training pipelines tailored for these smaller architectures. Through extensive experimental validation and analysis, our work aims to provide actionable recommendations for maximizing the reasoning capabilities of 0.5B models.
@@ 1 Introduction
The field of Language Models (LMs) has experienced a significant transformation with the advent of increasingly large models, which have demonstrated outstanding performance across a range of tasks [1, 2, 3]. However, the pursuit of such massive models has introduced considerable challenges, including substantial computational requirements, increased energy consumption, and growing privacy concerns. In this context, Small Reasoning Language Models (SRLMs) with approximately 0.5 billion parameters have emerged as a compelling alternative [4, 5]. These models not only provide enhanced computational efficiency but also offer cost-effectiveness, making them especially suitable for deployment in resource-constrained environments.
Despite their efficiency, SRLMs face inherent limitations due to their reduced size. The constrained capacity of 0.5 billion parameters presents a significant challenge when addressing complex tasks such as mathematical reasoning and code generation. These tasks require not only an understanding of intricate concepts but also the ability to produce precise and logical outputs. Early empirical evidence suggests that while SRLMs show potential in these areas, they exhibit a noticeable performance gap compared to their larger counterparts. This gap underscores the need for effective enhancement techniques to improve the capabilities of 0.5 billion parameter models. In recent studies, researchers have employed a variety of training strategies and pipelines to align the output of LMs with target preferences, such as upervised fine-tuning (SFT), knowledge distillation (KD), and reinforcement learning (RL) [6, 7]. However, in specific tasks, the training strategies of SRLMs have not been clearly explored, which makes it quite challenging to explore their capability boundaries and application fields.
@@
RQ3: What is the optimal combination of SFT, RL, and KD methods for enhancing the performance of 0.5B models in these domains? Here, we will investigate the synergistic effects of integrating various training techniques. By experimenting with diverse hybrid approaches, we aim to identify the most effective pipeline that harnesses the strengths of SFT, RL, and KD. This question seeks to reveal how the combination of these methodologies can lead to significant enhancements in the performance of 0.5B models, thereby offering a practical framework for optimizing their capabilities in mathematical reasoning and code generation tasks.
@@
  • We provide a thorough analysis of the capabilities of 0.5B language models in mathematical reasoning and code generation, highlighting their potential and limitations.
@@ 2.1 Mathematical Reasoning
Mathematical reasoning is a complex cognitive task that requires understanding, manipulating, and solving mathematical problems. For language models, tasks involving mathematical reasoning necessitate comprehending mathematical concepts, applying logical reasoning, and performing calculations to reach correct solutions. This includes solving word problems, proving theorems, and carrying out algebraic manipulations. The model must show an understanding of mathematical principles, follow multi-step reasoning processes, and generate accurate, well-structured answers.
@@ 2.2 Small Reasoning Language Models
The field of language modeling has experienced remarkable evolution over the past decade, marked by significant advancements in both methodology and application. Early language models, characterized by their relatively small size and limited capacity, struggled to capture the complexities inherent in human language. These initial models were often constrained by limited datasets, resulting in suboptimal performance across a diverse range of linguistic tasks. The landscape began to transform with the introduction of pre-training techniques, where researchers found that training on extensive corpora of text led to considerable enhancements in models’ abilities to comprehend and generate natural language [1, 2, 3]. This pivotal shift heralded the era of large language models (LLMs), which rapidly became the predominant paradigm in natural language processing.
The emergence of LLMs represented a significant milestone in AI research, epitomized by models such as GPT-3 [8], which boasted billions of parameters and demonstrated unprecedented proficiencies in language understanding and generation. These models exhibited remarkable capabilities across a wide array of tasks, including text completion, translation, question answering, and summarization. Their success was largely attributed to their immense size, enabling them to discern intricate patterns and relationships within vast datasets. However, the substantial computational and resource demands associated with LLMs posed significant challenges [9]; training and deploying these models required extensive computational power, considerable energy consumption, and substantial financial investment, thereby limiting accessibility for smaller research teams and organizations with constrained resources.
In light of the limitations presented by LLMs, researchers began to explore the potential of developing smaller models that could deliver comparable performance while mitigating resource requirements. This exploration led to the advent of Small Reasoning Language Models (SRLMs), typically defined as models possessing approximately 0.5-1.0 billion parameters, such as TinyLlama [10], Qwen2.5 series [11] and Qwen3 series [12]. SRLMs aim to strike a balance between performance and efficiency, rendering them suitable for deployment in resource-constrained environments. The impetus for SRLMs development arose from several key factors. Firstly, there was an escalating demand for models capable of operating on edge devices and within environments characterized by limited computational resources [5]. Secondly, privacy concerns associated with cloud-based LLMs prompted a search for models that could be executed locally without jeopardizing sensitive data. Finally, the environmental impact of training and deploying large models became a pressing concern, making smaller models an attractive and sustainable alternative.
SRLMs have found diverse applications across numerous domains where resource efficiency and deployment flexibility are vital. Their compact size renders them ideal for utilization in mobile applications, Internet of Things devices, and other contexts with limited computational capacity [13]. Furthermore, their reduced computational requirements facilitate faster response times and lower latency, which are critical for real-time applications. In the realm of mathematical reasoning, SRLMs have been successfully employed to develop systems capable of solving mathematical problems, assisting in educational contexts, and supporting research endeavors in mathematics. In the area of code generation, they have been instrumental in creating tools that enhance programmers’ productivity, automate routine programming tasks, and aid in the acquisition of new programming languages.
Despite their smaller size, SRLMs have demonstrated commendable performance across various tasks. Through rigorous optimization and the adoption of advanced training techniques, researchers have been able to enhance SRLMs capabilities, allowing them to approach the performance levels of their larger counterparts. This progress has unveiled new opportunities for applying AI in scenarios where resource constraints previously posed significant barriers. However, the smaller size of SRLMs also imposes inherent limitations on their ability to learn complex patterns and relationships within data. Such constraints may hinder their performance on tasks requiring deep understanding and sophisticated reasoning, particularly in advanced mathematical problem-solving [14] and intricate code generation [15, 16]. To address these challenges, researchers have developed a range of enhancement techniques, including supervised fine-tuning, reinforcement learning [7], and knowledge distillation [17]. These methodologies aim to maximize the potential of SRLMs while expanding their capabilities.
@@
The objective of this research is to investigate these enhancement techniques and their applications to 0.5 billion parameter models within the specific domains of mathematical reasoning and code generation. By tackling the challenges inherent in SRLMs and capitalizing on the opportunities they present, we aspire to contribute to the advancement of efficient, specialized AI models that deliver high performance while minimizing resource requirements.
@@ 3.1 Data Preparation
In this study, we utilize the GSM8K training set as a primary resource for both SFT and RL training intended to enhance the mathematical reasoning capabilities of 0.5B models. It comprises 7K multi-step arithmetic problems and undergoes a rigorous preprocessing regimen that includes deduplication, tokenization, and Chain of Thought (CoT) augmentation. This process is designed to create well-structured examples that facilitate reasoning. In addition to advancing mathematical reasoning, this training initiative aims to explore the potential for knowledge transfer. Specifically, we evaluate whether the logical and structured reasoning skills acquired through GSM8K can enhance the models’ performance in other math reasoning.
In this study, we utilize the GSM8K training set [18] as a primary resource for both SFT and RL training intended to enhance the mathematical reasoning capabilities of 0.5B models. It comprises 7K multi-step arithmetic problems and undergoes a rigorous preprocessing regimen that includes deduplication, tokenization, and Chain of Thought (CoT) augmentation [19]. This process is designed to create well-structured examples that facilitate reasoning. In addition to advancing mathematical reasoning, this training initiative aims to explore the potential for knowledge transfer. Specifically, we evaluate whether the logical and structured reasoning skills acquired through GSM8K can enhance the models’ performance in other math reasoning.
@@ 3.2 Foundation Models
We investigate the performance of three widely utilized SRLMs, including the Qwen2.5-0.5B-Instruct [11], Qwen2.5-0.5B [11], to assess their efficacy across various natural language processing tasks, particularly in mathematical reasoning and code generation. The Qwen2.5-0.5B-Instruct is specifically engineered for instruction-following tasks, demonstrating significant proficiency in generating responses that closely align with user prompts. In contrast, the Qwen2.5-0.5B serves as a versatile, general-purpose model, providing a robust foundation for a multitude of applications.
@@ 3.3 Training Strategies for Exploration
In this study, we employ the Group Relative Policy Optimization (GRPO) algorithm in reinforcement learning (RL), as it mitigates the necessity for a value model, reduces memory requirements, and allows for a clear articulation of reward targets, making it an optimal choice for the effective optimization of the policy model [20]. A pivotal aspect of GRPO is the design of an effective reward function, which must provide clear and informative signals to guide the model’s learning process. Consequently, we have developed a specific reward function tailored to enhance the model’s performance in mathematical reasoning and code generation tasks. The formulation of the reward function is delineated as follows:
@@
Knowledge Distillation (KD) transfers capabilities from large teacher models to smaller student models. In our pursuit of effective knowledge distillation for SRLMs, we employ the GSM8K distillation dataset, which is generated using the CAMEL framework [21]. This dataset comprises a collection of mathematical problem-solving traces, where each entry includes a problem statement accompanied by a comprehensive step-by-step solution. Our adherence to the methodology outlined by the CAMEL framework ensures that the distillation data is of high quality and maintains a structured format, which is essential for the efficient transfer of knowledge from the teacher model to our smaller models. By leveraging this specialized dataset, we aim to significantly enhance the mathematical reasoning capabilities of our models through the process of knowledge distillation.
@@ 3.4 Evaluation Benchmark & Metric
The evaluation of SRLMs training is conducted across several benchmarks, including OlympiadBench [22], MATH500 [23], MINERVA [24], AMC23 [25], and GSM8K test set [18]. OlympiadBench consists of advanced mathematical problems similar to those found in mathematical olympiads. MATH500 challenges models with complex mathematical concepts and problem-solving techniques. MINERVA focuses on scientific reasoning, requiring the integration of scientific knowledge with mathematical skills. AMC23 offers a unique set of mathematical challenges, while GSM8K specifically assesses the ability to handle grade school-level arithmetic problems that involve multi-step reasoning. Each benchmark is selected for its unique contribution to evaluating different aspects of mathematical reasoning.
Performance of each LMs will be measured using accuracy as the primary metric across all mathematical reasoning benchmarks . All evaluations are conducted using the LightEval framework [26] with the following configuration: models are evaluated with a maximum model length of 32768 tokens and 95% GPU memory utilization, using bfloat16 data type and data parallel processing across multiple GPUs. Generation parameters include maximum new tokens of 32768, temperature of 0.6, and top-p sampling of 0.95 to ensure robust and consistent evaluation results. We will also report computational costs, including training time and GPU resources utilized for each enhancement pipeline.
3.5 Implementation Settings
3.6 Environment
Table 1: Benchmarking LLMs at different scales. We follow the results of 7B and 1.5B models from [27, 28]. represents that we can not extract answer from model’s response. Dash (-) represents unavailable official results. All results are evaluated by Lighteval framework [26].
Foundation Model OlympiadBench MATH500 MINERVA AMC23 GSM8K
@@ 3.6 Environment
4 Experimental Results
4.1 Experiment 1: Comparison and Analysis of Different LMs
As demonstrated in Table 1, there is a distinct performance disparity between SRLMs and the larger scale LMs across various mathematical reasoning benchmarks. Specifically, the Qwen2.5-0.5B-Instruct model demonstrates a baseline performance of merely 6.2% on OlympiadBench and 31.4% on MATH500, significantly lagging behind the 7 billion and 1.5 billion parameter models, while the Qwen2.5-7B-Instruct model achieves an impressive 38.2% on OlympiadBench and 75.2% on MATH500, Qwen2.5-1.5B-Instruct model attains 16.7% on OlympiadBench and 51.2% on MATH500. These results underscore the inherent limitations associated with 0.5B SRLMs, attributing their subpar performance to their Relatively small parameter scale and week reasoning ability.
The implementation of effective training strategies has led to significant enhancements in the performance of 0.5B models. For instance, the Qwen2.5-0.5B-Instruct model, following optimization, demonstrates accuracy of 7.6% on OlympiadBench and 32.4% on MATH500. Despite significant enhancements, the performance of 0.5B SRLMs continues to fall short compared to their larger LMs. However, some optimized SRLMs have demonstrated a notable reduction in this performance gap. For example, Qwen3-0.6B [12] achieve performance metrics of 20.2% on OlympiadBench and 56.4% on MATH500 after undergoing similar training optimizations. These results underscore that appropriate training methodologies can effectively extend the capability boundaries of SRLMs. With further refinements in training strategies and improvements in model architecture, 0.5B models have the potential to attain performance levels that are increasingly comparable to those of larger models. This outcome demonstrates that the true capability boundaries of SRLMs are not static; rather, they can be expanded through optimized training methodologies.
Table 2: Comparison of different training strategies. represents that we can not extract answer from model’s response. Dash (-) represents unavailable official results. All results are evaluated by Lighteval framework [26].
Method OlympiadBench MATH500 MINERVA AMC23 GSM8K
@@ 4.3 Experiment 3: Comparison and Analysis of Hybrid Training Strategies
Table 3: Comparison of different hybrid training strategies.  represents that we can not extract answer from model’s response.  represents that the model fails to converge the reward curve during RL training, resulting in training collapse. Dash (-) represents unavailable official results. All results are evaluated by Lighteval framework [26].
Method OlympiadBench MATH500 MINERVA AMC23 GSM8K

### v3

# Effective Learning for Small Reasoning Models: An Empirical Study on 0.5B Reasoning LLMs

@@ Abstract
The ongoing evolution of language models has led to the development of large-scale architectures that demonstrate exceptional performance across a wide range of tasks. However, these models come with significant computational and energy demands, as well as potential privacy implications. In this context, Small Reasoning Language Models (SRLMs) with approximately 0.5 billion parameters present a compelling alternative due to their remarkable computational efficiency and cost-effectiveness, particularly in resource-constrained environments. Despite these advantages, the limited capacity of 0.5 billion parameter models poses challenges in handling complex tasks such as mathematical reasoning. This research investigates various training strategies, including supervised fine-tuning (SFT), knowledge distillation (KD), and reinforcement learning (RL), as well as their hybrid implementations, to enhance the performance of 0.5B SRLMs. We analyze effective methodologies to bridge the performance gap between SRLMS and larger models and present insights into optimal training pipelines tailored for these smaller architectures. Through extensive experimental validation and analysis, our work aims to provide actionable recommendations for maximizing the reasoning capabilities of 0.5B models.
@@ 1 Introduction
The field of Language Models (LMs) has experienced a significant transformation with the advent of increasingly large models, which have demonstrated outstanding performance across a range of tasks [zhao2023survey, liu2024survey, hadi2023survey]. However, the pursuit of such massive models has introduced considerable challenges, including substantial computational requirements, increased energy consumption, and growing privacy concerns. In this context, Small Reasoning Language Models (SRLMs) with approximately 0.5 billion parameters have emerged as a compelling alternative [van2024survey, lu2024small]. These models not only provide enhanced computational efficiency but also offer cost-effectiveness, making them especially suitable for deployment in resource-constrained environments.
Despite their efficiency, SRLMs face inherent limitations due to their reduced size. The constrained capacity of 0.5 billion parameters presents a significant challenge when addressing complex tasks such as mathematical reasoning. These tasks require not only an understanding of intricate concepts but also the ability to produce precise and logical outputs. Early empirical evidence suggests that while SRLMs show potential in these areas, they exhibit a noticeable performance gap compared to their larger counterparts. This gap underscores the need for effective enhancement techniques to improve the capabilities of 0.5 billion parameter models. In recent studies, researchers have employed a variety of training strategies and pipelines to align the output of LMs with target preferences, such as supervised fine-tuning (SFT), knowledge distillation (KD), and reinforcement learning (RL) [guo2025deepseek, dang2025reinforcement]. However, in specific tasks, the training strategies of SRLMs have not been clearly explored, which makes it quite challenging to explore their capability boundaries and application fields.
@@
RQ3: What is the optimal combination of SFT, RL, and KD methods for enhancing the performance of 0.5B models in these domains? Here, we will investigate the synergistic effects of integrating various training techniques. By experimenting with diverse hybrid approaches, we aim to identify the most effective pipeline that harnesses the strengths of SFT, RL, and KD. This question seeks to reveal how the combination of these methodologies can lead to significant enhancements in the performance of 0.5B models, thereby offering a practical framework for optimizing their capabilities in mathematical reasoning tasks.
@@
  • We provide a thorough analysis of the capabilities of 0.5B language models in mathematical reasoning, highlighting their potential and limitations.
@@ 2.1 Mathematical Reasoning
Mathematical reasoning is a complex cognitive task that requires understanding, manipulating, and solving mathematical problems [Creswell2022, Drori2022, Lewkowycz2022]. For language models, tasks involving mathematical reasoning necessitate comprehending mathematical concepts, applying logical reasoning, and performing calculations to reach correct solutions [Cobbe2021, Lample2022, Frieder2023]. This includes solving word problems, proving theorems, and carrying out algebraic manipulations [Hendrycks2021]. The model must show an understanding of mathematical principles, follow multi-step reasoning processes, and generate accurate, well-structured answers [Wang2023, Gao2022, liu2023mathematical, magister2023teaching, ahn2024large, sun2024mm, lu2023mathvista].
@@ 2.2 Small Reasoning Language Models
Over the past decade, language modeling has evolved significantly, transitioning from early models limited by size and data to the era of large language models (LLMs), which leverage pre-training on extensive corpora to achieve remarkable natural language understanding and generation [zhao2023survey, liu2024survey, hadi2023survey]. LLMs, exemplified by models like GPT-3 [floridi2020gpt], possess billions of parameters and excel at tasks such as text completion, translation, and summarization, yet their immense computational demands [kaddour2023challenges] restrict accessibility for smaller organizations. In response, researchers have developed Small Reasoning Language Models (SRLMs)—typically with 0.5-1.0 billion parameters, such as TinyLlama [zhang2024tinyllama], Qwen2.5 [qwen2.5], and Qwen3 [qwen3]—which balance performance and efficiency for deployment in resource-constrained environments. Motivated by the need for edge-device compatibility [lu2024small], privacy, and sustainability, SRLMs offer practical solutions for mobile and IoT applications [li2025efficient], with benefits including lower latency and faster response times. Their use extends to mathematical reasoning, education, and code generation, supporting both productivity and learning. Although SRLMs achieve impressive results through advanced optimization and training, their smaller size imposes limitations on complex reasoning tasks [ranaldi2024aligning, souza2025code, sqlr1]. To address these challenges, enhancement techniques such as supervised fine-tuning, reinforcement learning [dang2025reinforcement], and knowledge distillation [chen2025sft] have been developed, further expanding their capabilities while maintaining resource efficiency.
@@
The objective of this research is to investigate these enhancement techniques and their applications to 0.5 billion parameter models within the specific domains of mathematical reasoning. By tackling the challenges inherent in SRLMs and capitalizing on the opportunities they present, we aspire to contribute to the advancement of efficient, specialized AI models that deliver high performance while minimizing resource requirements.
@@ 3.1 Data Preparation
In this study, we utilize the GSM8K training set [cobbe2021gsm8k] as a primary resource for both SFT and RL training intended to enhance the mathematical reasoning capabilities of 0.5B models. It comprises 7K multi-step arithmetic problems and undergoes a rigorous preprocessing regimen that includes deduplication, tokenization, and Chain of Thought (CoT) augmentation [wei2022chain]. This process is designed to create well-structured examples that facilitate reasoning. In addition to advancing mathematical reasoning, this training initiative aims to explore the potential for knowledge transfer. Specifically, we evaluate whether the logical and structured reasoning skills acquired through GSM8K can enhance the models’ performance in other math reasoning.
@@ 3.2 Foundation Models
We investigate the performance of three widely utilized SRLMs, including the Qwen2.5-0.5B-Instruct [qwen2.5], Qwen2.5-0.5B [qwen2.5], to assess their efficacy across various natural language processing tasks, particularly in mathematical reasoning. The Qwen2.5-0.5B-Instruct is specifically engineered for instruction-following tasks, demonstrating significant proficiency in generating responses that closely align with user prompts. In contrast, the Qwen2.5-0.5B serves as a versatile, general-purpose model, providing a robust foundation for a multitude of applications.
@@ 3.3 Training Strategies for Exploration
In this study, we employ the Group Relative Policy Optimization (GRPO) algorithm in reinforcement learning (RL), as it mitigates the necessity for a value model, reduces memory requirements, and allows for a clear articulation of reward targets, making it an optimal choice for the effective optimization of the policy model [grpo]. A pivotal aspect of GRPO is the design of an effective reward function, which must provide clear and informative signals to guide the model’s learning process. Consequently, we have developed a specific reward function tailored to enhance the model’s performance in mathematical reasoning tasks. The formulation of the reward function is delineated as follows:
@@
Knowledge Distillation (KD) transfers capabilities from large teacher models to smaller student models. In our pursuit of effective knowledge distillation for SRLMs, we employ the GSM8K distillation dataset, which is generated using the CAMEL framework [li2023camel]. This dataset comprises a collection of mathematical problem-solving traces, where each entry includes a problem statement accompanied by a comprehensive step-by-step solution. Our adherence to the methodology outlined by the CAMEL framework ensures that the distillation data is of high quality and maintains a structured format, which is essential for the efficient transfer of knowledge from the teacher model to our smaller models. By leveraging this specialized dataset, we aim to significantly enhance the mathematical reasoning capabilities of our models through the process of knowledge distillation.
@@ 3.4 Evaluation Benchmark & Metric
The evaluation of SRLMs training is conducted across several benchmarks, including OlympiadBench [he2024olympiadbench], MATH500 [hendrycks2021measuring], MINERVA [Minervamath], AMC23 [AMC23], and GSM8K test set [cobbe2021gsm8k]. OlympiadBench consists of advanced mathematical problems similar to those found in mathematical olympiads. MATH500 challenges models with complex mathematical concepts and problem-solving techniques. MINERVA focuses on scientific reasoning, requiring the integration of scientific knowledge with mathematical skills. AMC23 offers a unique set of mathematical challenges, while GSM8K specifically assesses the ability to handle grade school-level arithmetic problems that involve multi-step reasoning. Each benchmark is selected for its unique contribution to evaluating different aspects of mathematical reasoning.
Performance of each LMs will be measured using accuracy as the primary metric across all mathematical reasoning benchmarks . All evaluations are conducted using the LightEval framework [lighteval] with the following configuration: models are evaluated with a maximum model length of 32768 tokens and 95% GPU memory utilization, using bfloat16 data type and data parallel processing across multiple GPUs. Generation parameters include maximum new tokens of 32768, temperature of 0.6, and top-p sampling of 0.95 to ensure robust and consistent evaluation results. We will also report computational costs, including training time and GPU resources utilized for each enhancement pipeline.
4 Experimental Results
4.1 Experiment 1: Comparison and Analysis of Different LMs
As demonstrated in Table 1, there is a distinct performance disparity between SRLMs and the larger scale LMs across various mathematical reasoning benchmarks. Specifically, the Qwen2.5-0.5B-Instruct model demonstrates a baseline performance of merely 6.2% on OlympiadBench and 31.4% on MATH500, significantly lagging behind the 7 billion and 1.5 billion parameter models, while the Qwen2.5-7B-Instruct model achieves an impressive 38.2% on OlympiadBench and 75.2% on MATH500, Qwen2.5-1.5B-Instruct model attains 16.7% on OlympiadBench and 51.2% on MATH500. These results underscore the inherent limitations associated with 0.5B SRLMs, attributing their subpar performance to their relatively small parameter scale and weak reasoning ability.
The implementation of effective training strategies has led to significant enhancements in the performance of 0.5B models. For instance, the Qwen2.5-0.5B-Instruct model, following optimization, demonstrates accuracy of 7.6% on OlympiadBench and 32.4% on MATH500. Despite significant enhancements, the performance of 0.5B SRLMs continues to fall short compared to their larger LMs. However, some optimized SRLMs have demonstrated a notable reduction in this performance gap. For example, Qwen3-0.6B [qwen3] achieve performance metrics of 20.2% on OlympiadBench and 56.4% on MATH500 after undergoing similar training optimizations. These results underscore that appropriate training methodologies can effectively extend the capability boundaries of SRLMs. With further refinements in training strategies and improvements in model architecture, 0.5B models have the potential to attain performance levels that are increasingly comparable to those of larger models. This outcome demonstrates that the true capability boundaries of SRLMs are not static; rather, they can be expanded through optimized training methodologies.
Table 1: Benchmarking LLMs at different scales. We follow the results of 7B and 1.5B models from [deepscaler2025, zeng7b]. OB represents OlympiadBench. represents that we can not extract answer from model’s response. Dash (-) represents unavailable official results. All results are evaluated by Lighteval framework [lighteval].
Foundation Model OB MATH500 MINERVA AMC23 GSM8K
@@ 4.1 Experiment 1: Comparison and Analysis of Different LMs
Table 2: Comparison of different training strategies. OB represents OlympiadBench. represents that we can not extract answer from model’s response. Dash (-) represents unavailable official results. All results are evaluated by Lighteval framework [lighteval].
Method OB MATH500 MINERVA AMC23 GSM8K
@@
Table 3: Comparison of different hybrid training strategies. OB represents OlympiadBench.  represents that we can not extract answer from model’s response.  represents that the model fails to converge the reward curve during RL training, resulting in training collapse. Dash (-) represents unavailable official results. All results are evaluated by Lighteval framework [lighteval].
Method OB MATH500 MINERVA AMC23 GSM8K